# Model-Agnostic, Temperature-Informed Sampling Enhances Cross-Year Crop Mapping with Deep Learning

### v2

@@ 3.3 T3S: Thermal-Time-based Temporal Sampling
Satellite-based crop monitoring approaches often rely on fixed-interval calendar-day sampling to construct input sequences. However, this overlooks the ecophysiological processes of plant growth such as their dependence on temperature. As crop development is tightly coupled with temperature, we propose a simple yet effective model-agnostic sampling method, Thermal-Time-based Temporal Sampling (T3S), which replaces fixed calendar intervals with intervals defined in thermal time, using Growing Degree Days (GDD) as the temporal reference.
GDD provides a cumulative measure of heat accumulation, reflecting the physiological development of crops more faithfully than absolute time. For a given day $d$, GDD is computed as:
$\mathrm{GDD}_{d}=\max\left(0,\frac{T_{\text{max},d}+T_{\text{min},d}}{2}-T_{\text{base}}\right),$ (1)
where $T_{\text{max},d}$ and $T_{\text{min},d}$ are the daily maximum and minimum temperatures, respectively, and $T_{\text{base}}$ is a base threshold below which crop development is considered negligible. In our experiments, we use $T_{\text{base}}=0^{\circ}$, a commonly used value for temperate climate crops.
We then compute the cumulative GDD over time:
$\mathrm{GDD}_{D}=\sum_{d=1}^{D}\mathrm{GDD}_{d}.$ (2)
To construct a time series of $L$ observations, we partition the full GDD range $\left[\mathrm{GDD}{\min},\mathrm{GDD}{\max}\right]$ into $L$ uniform intervals. Within each interval, we select the least cloudy satellite observation, prioritizing data quality. Importantly, Sentinel-2 products include a freely available cloud mask layer, allowing direct use of cloud information without additional preprocessing. This ensures that the selected timestamps correspond to ecophysiological meaningful intervals that are consistent across different years, even when calendar dates diverge significantly due to seasonal variability.
The T3S algorithm is outlined in Algorithm 1. It takes as input the full set of timestamps, the cumulative thermal time values, i.e, GDD, the satellite data cube, a cloud mask for each observation, and the desired number of samples $L$. For each thermal interval, it identifies candidate observations and selects the one with the fewest cloud-contaminated pixels. The final set of selected indices is sorted to preserve temporal order. Figure 5 illustrates the overall workflow of our approach.
@@
Figure 5: The first row shows conventional uniform sampling; the second row shows uniform sampling where the least cloudy observation is chosen in each interval (referred to as “Deformable” baseline); the third row shows the proposed Thermal‐Time‐based Temporal Sampling (T3S) approach. In T3S, each square box represents the temperature sum accumulated since the previous one, illustrating intervals with equal thermal time.
@@
1:function T3S(timeStamps, GDD, data, cloudMask, $L$)
2:     $t_{\min}\leftarrow\min(\text{GDD})$
3:     $t_{\max}\leftarrow\max(\text{GDD})$
4:     $\Delta\leftarrow(\,t_{\max}-t_{\min}\,)/L$   #$L$ is target temporal length
@@
7:     while $|selected|<L$ and $current<t_{\max}$ do
@@
9:         $window\leftarrow\{\,i\mid current\leq\text{GDD}[i]<next\}$
@@
11:              $best\leftarrow\displaystyle\arg\max_{i\in window}\sum(\text{cloudMask}[i]=0)$
@@
17:     return data[:, selected, …]
@@
3.4 Training Details
The model was trained for 100 epochs using a batch size of 8 and a fixed temporal length of 24. We employed the Adam optimizer with an initial learning rate of 0.001, scheduled by a OneCycleLR policy with a maximum learning rate of 0.01 and a warm-up phase covering 30% of the total training steps. The U-TAE architecture is configured with an encoder having channel widths $[128,128,128,256]$, a symmetric decoder with widths $[32,64,128,256]$ and strided convolutions with kernel size $k=4$, stride $s=2$, and padding $p=1$. The temporal attention module employs $n_{\text{head}}=16$ heads and a model dimension of $d_{\text{model}}=256$. Input features are standardized to zero mean and unit variance on a per-channel basis. We implemented the model in PyTorch and trained it on a single NVIDIA RTX 4090 GPU over roughly three days. The full codebase and pretrained checkpoints will be publicly available on GitHub.
@@ 4 Dataset
In this work, we work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland from 2021 to 2023. SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery with annual crop‐type labels for each field. The dataset includes 50 distinct crop types, each representing the primary crop grown in a field during a given season. These labels where provided by farmers to the cantons of Switzerland according to a data model provided by the Federal Office for Agriculture  [^1] https://www.blw.admin.ch/de/landwirtschaftliche-kulturflaechen and joined together by the Swiss Centre for Agricultural Research, Agroscope. Unlike many existing datasets, e.g. Pastis dataset (Garnot and Landrieu,, 2021), SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (cf. Figure 6).
[…]
To complement these remote sensing observations, we incorporate high-resolution gridded climate data from MeteoSwiss, which supplies near-surface temperature fields (minimum, maximum, and average) on 1 km, 2 km, and 5 km grids (Federal Office of Meteorology and Climatology, MeteoSwiss,, 2024). In our study, we use the 1 km daily minimum (TminD) and maximum (TmaxD) temperature series to compute growing degree days (GDD; refer to Section 3.3).
5 Experiment & Results
We evaluate the proposed T3S against four very relevant state-of-the-art baselines using a leave‐one‐year‐out scheme over the 2021–2023 SwissCrop dataset. In each of six cross‐year folds, two seasons are held out for testing while the remaining one season is used for training. To further assess robustness, we conduct a low‐data experiment using only 10% of the labels, and an early‐season study by truncating each test timeseries at the 50th and 75th time percentile (end of June and end of September respectively). We assess model predictive performance using Overall accuracy, the fraction of pixels correctly labeled; Intersection‐over‐Union (IoU), the ratio of the area of overlap between predicted and ground‐truth pixels to the area of their union; and Mean Intersection‐over‐Union (mIoU), averaged across classes. Calibration is measured by the Expected Calibration Error (ECE), which measures how far predicted confidences deviate from observed error rates, i.e., perfect calibration occurs when the estimated uncertainties match the actual likelihood of a misclassification. We also report Negative Log‐Likelihood (NLL), the mean negative log‐probability of the true class, and the Brier Score, the mean squared error between predicted probabilities and one‐shot labels.
@@ 5 Experiment & Results
2021 2022 U-TAE (Garnot and Landrieu,, 2021) 70.1 17.2 53.9 0.053 0.943 0.412
+MC-Dropout (Gal and Ghahramani,, 2016) 71.3 17.4 55.4 0.032 0.911 0.402
+Thermal-PE (Nyborg et al.,, 2022) 72.5 18.6 56.9 0.045 0.854 0.385
@@
2021 2023 U-TAE (Garnot and Landrieu,, 2021) 71.4 16.7 55.6 0.052 0.947 0.399
+MC-Dropout (Gal and Ghahramani,, 2016) 72.5 17.0 56.8 0.037 0.918 0.394
+Thermal-PE (Nyborg et al.,, 2022) 71.3 15.5 55.4 0.035 0.925 0.401
@@
2022 2021 U-TAE (Garnot and Landrieu,, 2021) 70.8 16.2 54.8 0.035 0.936 0.408
+MC-Dropout (Gal and Ghahramani,, 2016) 71.2 16.2 55.3 0.024 0.913 0.399
+Thermal-PE (Nyborg et al.,, 2022) 72.8 18.2 57.2 0.077 1.022 0.403
@@
2022 2023 U-TAE (Garnot and Landrieu,, 2021) 72.1 17.4 56.4 0.041 0.875 0.389
+MC-Dropout (Gal and Ghahramani,, 2016) 72.8 17.2 57.2 0.028 0.853 0.381
+Thermal-PE (Nyborg et al.,, 2022) 63.9 14.2 46.9 0.123 1.374 0.525
@@
2023 2021 U-TAE (Garnot and Landrieu,, 2021) 71.6 17.6 55.8 0.066 0.913 0.398
+MC-Dropout (Gal and Ghahramani,, 2016) 71.6 17.6 55.7 0.040 0.889 0.391
+Thermal-PE (Nyborg et al.,, 2022) 71.8 16.3 56.0 0.098 0.987 0.413
@@
2023 2022 U-TAE (Garnot and Landrieu,, 2021) 71.2 18.4 55.3 0.059 0.957 0.410
+MC-Dropout (Gal and Ghahramani,, 2016) 72.0 18.6 56.2 0.043 0.929 0.400
+Thermal-PE (Nyborg et al.,, 2022) 75.3 19.6 60.4 0.085 0.929 0.366
@@
Average U-TAE (Garnot and Landrieu,, 2021) $71.2\pm 0.6$ $17.3\pm 0.7$ $55.3\pm 0.8$ $0.051\pm 0.012$ $0.929\pm 0.028$ $0.403\pm 0.008$
+MC-Dropout (Gal and Ghahramani,, 2016) $71.9\pm 0.6$ $17.3\pm 0.7$ $56.1\pm 0.7$ $\underline{0.034}\pm 0.007$ $0.902\pm 0.029$ $0.395\pm 0.007$
+Thermal-PE (Nyborg et al.,, 2022) $71.3\pm 3.5$ $17.1\pm 1.9$ $55.5\pm 4.1$ $0.077\pm 0.029$ $1.015\pm 0.156$ $0.416\pm 0.053$
@@
Figure 8: Class-wise accuracy difference between the proposed T3S and U-TAE. Averaged over six experimental folds. Blue bars indicate classes where T3S outperforms U-TAE, while red bars indicate classes where T3S underperforms.
@@
U-TAE is a U‐Net combined with a temporal attention encoder (i.e., self-attention layer) that ingests images at fixed calendar dates and uses sinusoidal day‐of‐year positional encodings; it serves as our primary reference. We combine U-TAE with MC Dropout (Gal and Ghahramani,, 2016): U-TAE+MC‐Dropout retains the same sampling grid but activates dropout layers at test time to perform Monte Carlo sampling, thereby approximating a Bayesian posterior and yielding pixel‐level uncertainty estimates without changing the input dates. Note that Monte Carlo Dropout remains one of the most popular and theoretically grounded methods for uncertainty calibration in deep learning (Kendall and Gal,, 2017). Empirically, the number of ensemble members is set to 5, and the drop rate is set to 0.2. U-TAE+Thermal-PE follows Nyborg et al., (2022) and replaces the sinusoidal calendrical encodings with absolute thermal‐time (Growing Degree Days) embeddings, giving the network explicit phenological context while still sampling on calendar dates. U-TAE+Deformable Sampling follows the same fixed‐interval calendar schedule (e.g. one image per fifteen days) but “deforms” each interval by selecting the least‐cloudy observation, thus reducing noise from clouds without altering the temporal grid.
@@
Crucially, T3S also yields much better‐calibrated outputs. T3S reduces ECE by roughly fourfold compared to U-TAE, approaching perfect calibration, and substantially lowers both Negative Log‐Likelihood and Brier Score, indicating sharper and more reliable uncertainty estimates. By comparison, U-TAE+MC‐Dropout achieves the second‐best calibration, modestly boosting accuracy and reducing ECE relative to vanilla U-TAE but still falling short of T3S. Note that MC-Dropout is specifically designed to improve the uncertainty calibration. U-TAE+Thermal-PE introduces phenological context via GDD embeddings and yields moderate gains in both performance and calibration, yet does not match the robustness of thermal‐time sampling. Refer to Section 3.2 for a detailed analysis of its limited efficacy. Deformable Sampling improves predictive accuracy through cloud filtering but does so at the expense of uncertainty quality, exhibiting higher ECE. Overall, these results highlight that aligning observations by thermal time via temporal sampling, combined with cloud‐aware selection yields the most robust and trustworthy cross‐year crop mapping models.
@@
Figure 11 contrasts T3S and U-TAE on both classification and uncertainty estimation, where pixel-wise uncertainty is defined as $1-\max(\mathrm{softmax})\in[0,1]$. In the first example, both models misclassify the highlighted region, but U-TAE remains overconfident in its incorrect label while T3S appropriately elevates uncertainty around the error. In the second example, T3S confines its errors to the fuzzy edges of the highlighted field, which is understandable given that farmers hand-draw plot boundaries and 10m pixels often cover two crop types, and its uncertainty map shows a pronounced hotspot exactly along those misclassified borders. In contrast, U-TAE not only mislabels the entire field but also produces a diffuse, misaligned uncertainty map that neither highlights its own errors nor respects the actual field geometry. In the third example, T3S again flags its error with a pronounced uncertainty hotspot, even revealing a likely ground-truth annotation mistake, while U-TAE provides no meaningful uncertainty signal. These examples illustrate that T3S not only reduces classification errors but also delivers well-calibrated uncertainty estimates that expose both model and dataset inconsistencies.
@@
In many agricultural settings, annotated crop labels are scarce or expensive to obtain, which makes low data performance a critical requirement for operational viability. To assess robustness under label scarcity, we conducted a six-fold experiment using only 10 % of the training data, randomly sampled once and held constant across methods, and evaluated both T3S and U-TAE on the full test set. As shown in Figure 9, both models incur drops in accuracy and increases in Expected Calibration Error when labels are limited, but U-TAE’s performance deteriorates far more rapidly. Remarkably, T3S trained on just 10 % of the data still outperforms U-TAE trained on the full dataset, and its calibration remains substantially more reliable.
This advantage stems from the incorporation of a meaningful inductive bias, namely the thermal time prior, which encodes domain knowledge about crop development into the model architecture. By constraining the hypothesis space to reflect biologically plausible growth patterns, T3S requires fewer examples to learn robust representations, making it far more data efficient and noise resilient, for example, against cloud cover and sensor artifacts, than models trained purely in a data-driven fashion. These results demonstrate that thermal time inductive bias confers strong resilience to small data scenarios, underscoring T3S’s practicality for real-world applications where labels are limited. See Section 6 for more discussion.
Early‐Season Classification.
Early season classification is an important aspect of this technology, since some applications require crop labeling during the growing season rather than after harvest. To evaluate T3S in this context, we truncate the test set time series at two percentiles, 75 percent (end of September) and 50 percent (end of June), and perform a six fold experiment in which both T3S and U-TAE are trained on full-season data and tested on the truncated series. Figure 10 shows that both models experience a decline in in-season performance, measured by predictive accuracy and uncertainty estimation, when using partial time series. At the 75 percent cutoff, U-TAE’s performance drops sharply, whereas T3S remains virtually unchanged. Moreover, at the 50 percent cutoff, T3S attains a similar level of accuracy and uncertainty calibration that U-TAE achieves using the full season of data. These results indicate that T3S not only improves in-season classification accuracy and uncertainty quantification but also underscores its practical suitability for operational in-season crop monitoring and decision-support applications, where timely and reliable predictions are critical.

### v3

@@ 3.3 T3S: Thermal-Time-based Temporal Sampling
Satellite-based crop monitoring often relies on fixed-interval calendar‑day sampling to construct input sequences, overlooking temperature‑driven plant ecophysiology. Since crop development is tightly coupled with accumulated heat, we introduce a simple yet effective, model‑agnostic sampling method, T3S, which reparametrizes calendar time into thermal time, replacing fixed calendar intervals with intervals defined by growing degree days (GDD) as the temporal reference, and then performs uniform subsampling in this thermal‑time domain.
GDD provides a cumulative measure of heat accumulation, reflecting the physiological development of crops more faithfully than absolute time. For a given day $i$, GDD is computed as
$\mathrm{GDD}_{i}=\max\left(0,\frac{T_{\text{max},i}+T_{\text{min},i}}{2}-T_{\text{base}}\right),$ (1)
where $T_{\text{max},i}$ and $T_{\text{min},i}$ are the daily maximum and minimum temperatures for day $i$, respectively, and $T_{\text{base}}$ is a base threshold below which crop development is considered negligible. In our experiments, we use $T_{\text{base}}=0^{\circ}$, a commonly used value for temperate climate crops. We then compute the cumulative GDD (cGDD) over time for day d:
$\mathrm{cGDD}_{d}=\sum_{i=1}^{d}\mathrm{GDD}_{i}.$ (2)
To construct a time series of $T(<L)$ observations, we partition the full cGDD range $\left[\mathrm{cGDD}{\min},\mathrm{cGDD}{\max}\right]$ into $T$ uniform intervals. Within each interval, we select the least cloudy satellite observation, prioritizing data quality. Importantly, Sentinel-2 products include a freely available cloud mask layer, allowing direct use of cloud information without additional preprocessing. This ensures that the selected timestamps correspond to ecophysiological meaningful intervals that are consistent across different years, even when calendar dates diverge significantly due to seasonal variability.
The T3S algorithm is outlined in Algorithm 1. It takes as input the full set of timestamps $\{1,2,...,L\}$, where $L$ is the initial time length, the satellite data $\bm{X}\in{\rm I\!R}^{L\times H\times W\times B}$, the cumulative thermal time values, i.e, $cGDD$, a cloud mask for each observation, and the desired number of samples $T$. For each thermal interval, it identifies candidate observations and selects the one with the fewest cloud-contaminated pixels. The final set of selected indices is sorted to preserve temporal order. Figure 5 illustrates the overall workflow of our approach.
@@
Figure 5: Method overview. The first row shows conventional uniform sampling; the second row shows uniform sampling where the least cloudy observation is chosen in each interval (referred to as “Deformable” baseline); the third row shows the proposed Thermal Time‐based Temporal Sampling (T3S). In T3S, each square box represents the temperature sum accumulated since the previous one, illustrating intervals with equal cumulative GDD.
@@
1:function T3S(timeStamps, $\bm{X}$, cGDD, cloudMask, $T$)
2:     $t_{\min}\leftarrow\min(\text{cGDD})$     # $\bm{X}$’s shape is [L,H,W,B]
3:     $t_{\max}\leftarrow\max(\text{cGDD})$
4:     $\Delta\leftarrow(\,t_{\max}-t_{\min}\,)/T$   #$T$ is target temp. length
@@
7:     while $|selected|<T$ and $current<t_{\max}$ do
@@
9:         $window\leftarrow\{\,i\mid current\leq\text{cGDD}[i]<next\}$
@@
11:              $best\leftarrow\displaystyle\operatorname*{arg\,max}_{i\in window}\sum(\text{cloudMask}[i]=0)$
@@
17:     return $\bm{X}$[selected, …]  # the shape is [T,H,W,B]
@@ 4 Dataset
We work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland (see Figure 10 in the appendix). SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery, averaging 110 timestamps per season, with annual crop‐type labels for each field. The dataset includes 50 distinct crop types, each representing the primary crop grown in a field during a given season. Unlike many existing datasets, e.g., commonly used Pastis benchmark [14], SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (see Figure 11 in the appendix for the class distribution of SwissCrop). To drive phenology-aware sampling, we complement the imagery with daily temperature data (originally at 1 km resolution and resampled to 10 m to match Sentinel-2 imagery) from MeteoSwiss to compute growing degree days (GDD). The full SwissCrop dataset, including imagery, labels, and temperature series, is publicly released to facilitate reproducible research in large-scale crop mapping, phenology modeling, and earth observation research. For more details, refer to the Appendix B.
5 Experiment & Discussion
We evaluate the proposed T3S against four relevant state-of-the-art baselines using a leave‐one‐year‐out scheme over the 2021–2023 SwissCrop dataset. In each of six cross‐year folds, two seasons are held out for testing while the remaining one season is used for training. To further assess robustness, we conduct a low‐data experiment using only 10% of the labels, and an early‐season study by truncating each test time series at the 50th and 75th time percentile (end of June and end of September, respectively). We assess model predictive performance using overall accuracy, the fraction of pixels correctly labeled; intersection‐over‐union (IoU), the ratio of the area of overlap between predicted and ground‐truth pixels to the area of their union; and mean intersection‐over‐union (mIoU), averaged across classes. Calibration is measured by the expected calibration error (ECE), which measures how far predicted confidences deviate from observed error rates, i.e., perfect calibration occurs when estimated uncertainties match the actual likelihood of misclassification. We also report negative log‐likelihood (NLL), the mean negative log‐probability of the true class, and Brier score, the mean squared error between predicted probabilities and one‐shot labels.
@@ 5 Experiment & Discussion
2021 2022 U-TAE [14] 70.1 17.2 53.9 0.053 0.943 0.412
+MC-Dropout [12] 71.3 17.4 55.4 0.032 0.911 0.402
+Thermal-PE [31] 72.5 18.6 56.9 0.045 0.854 0.385
@@
2021 2023 U-TAE [14] 71.4 16.7 55.6 0.052 0.947 0.399
+MC-Dropout [12] 72.5 17.0 56.8 0.037 0.918 0.394
+Thermal-PE [31] 71.3 15.5 55.4 0.035 0.925 0.401
@@
2022 2021 U-TAE [14] 70.8 16.2 54.8 0.035 0.936 0.408
+MC-Dropout [12] 71.2 16.2 55.3 0.024 0.913 0.399
+Thermal-PE [31] 72.8 18.2 57.2 0.077 1.022 0.403
@@
2022 2023 U-TAE [14] 72.1 17.4 56.4 0.041 0.875 0.389
+MC-Dropout [12] 72.8 17.2 57.2 0.028 0.853 0.381
+Thermal-PE [31] 63.9 14.2 46.9 0.123 1.374 0.525
@@
2023 2021 U-TAE [14] 71.6 17.6 55.8 0.066 0.913 0.398
+MC-Dropout [12] 71.6 17.6 55.7 0.040 0.889 0.391
+Thermal-PE [31] 71.8 16.3 56.0 0.098 0.987 0.413
@@
2023 2022 U-TAE [14] 71.2 18.4 55.3 0.059 0.957 0.410
+MC-Dropout [12] 72.0 18.6 56.2 0.043 0.929 0.400
+Thermal-PE [31] 75.3 19.6 60.4 0.085 0.929 0.366
@@
Average U-TAE [14] $71.2\pm 0.6$ $17.3\pm 0.7$ $55.3\pm 0.8$ $0.051\pm 0.012$ $0.929\pm 0.028$ $0.403\pm 0.008$
+MC-Dropout [12] $71.9\pm 0.6$ $17.3\pm 0.7$ $56.1\pm 0.7$ $\underline{0.034}\pm 0.007$ $0.902\pm 0.029$ $0.395\pm 0.007$
+Thermal-PE [31] $71.3\pm 3.5$ $17.1\pm 1.9$ $55.5\pm 4.1$ $0.077\pm 0.029$ $1.015\pm 0.156$ $0.416\pm 0.053$
@@
Figure 6: Class-wise accuracy difference between T3S and U-TAE. Averaged over six experimental folds. Blue bars indicate classes where T3S outperforms U-TAE, while red bars indicate classes where T3S underperforms.
@@
U-TAE is a U‐Net combined with a temporal attention encoder (i.e., self-attention layer) that ingests images at fixed calendar dates and uses sinusoidal day‐of‐year positional encodings; it serves as our primary reference. We combine U-TAE with MC Dropout [12]: U-TAE+MC‐Dropout retains the same sampling grid but activates dropout layers at test time to perform Monte Carlo sampling, thereby approximating a Bayesian posterior and yielding pixel‐level uncertainty estimates without changing the input dates. Note that Monte Carlo Dropout remains one of the most popular and theoretically grounded methods for uncertainty calibration in deep learning [21]. Empirically, the number of ensemble members is set to 5, and the drop rate is set to 0.2. U-TAE+Thermal-PE follows Nyborg et al. [31] and replaces the sinusoidal calendrical encodings with absolute thermal‐time (Growing Degree Days) embeddings, giving the network explicit phenological context while still sampling on calendar dates. U-TAE+Deformable Sampling follows the same fixed‐interval calendar schedule (e.g. one image per fifteen days) but “deforms” each interval by selecting the least‐cloudy observation, thus reducing noise from clouds without altering the temporal grid.
@@
Crucially, T3S also yields much better‐calibrated outputs. T3S reduces ECE by roughly fourfold compared to U-TAE, approaching perfect calibration, and substantially lowers both NLL and Brier score, indicating sharper and more reliable uncertainty estimates. By comparison, U-TAE+MC‐Dropout achieves the second‐best calibration, modestly boosting accuracy and reducing ECE relative to vanilla U-TAE but still falling short of T3S. Note that MC-Dropout is specifically designed to improve the uncertainty calibration. U-TAE+Thermal-PE introduces phenological context via GDD embeddings and yields moderate gains in both performance and calibration, yet does not match the robustness of thermal‐time sampling. Refer to Section 3.2 for a detailed analysis of its limited efficacy. Deformable Sampling improves predictive accuracy through cloud filtering but does so at the expense of uncertainty quality, exhibiting higher ECE. Overall, these results highlight that aligning observations by thermal time via temporal sampling, combined with cloud‐aware selection yields the most robust and trustworthy cross‐year crop mapping models.
@@
Figure 9 contrasts T3S and U-TAE on both classification and uncertainty estimation, where pixel-wise uncertainty is defined as $1-\max(\mathrm{softmax})\in[0,1]$. In the first example, both models misclassify the highlighted region, but U-TAE remains overconfident in its incorrect label while T3S appropriately elevates uncertainty around the error. In the second example, T3S confines its errors to the fuzzy edges of the highlighted field. This is understandable given that farmers hand-draw plot boundaries and 10m pixels often cover two crop types, and the T3S uncertainty map shows a pronounced hotspot exactly along those misclassified borders. In contrast, U-TAE not only mislabels the entire field but also produces a diffuse, misaligned uncertainty map that neither highlights its own errors nor respects the actual field geometry. In the third example, T3S again flags its error with a pronounced uncertainty hotspot, even revealing a likely ground-truth annotation mistake, while U-TAE provides no meaningful uncertainty signal. These examples illustrate that T3S not only reduces classification errors but also delivers well-calibrated uncertainty estimates that expose both model and dataset inconsistencies.
Early‐Season Classification.
Early season classification is an important aspect of this technology, since some applications require crop labeling during the growing season rather than after harvest. To evaluate T3S in this context, we truncate the test set time series at two percentiles, 75 percent (end of September) and 50 percent (end of June), and perform a six fold experiment in which both T3S and U-TAE are trained on full-season data and tested on the truncated series. Figure 7 shows that both models experience a decline in in-season performance when using partial time series. At the 75 percent cutoff, U-TAE’s performance drops sharply, whereas T3S remains virtually unchanged. Moreover, at the 50 percent cutoff, T3S attains a similar level of accuracy and uncertainty calibration that U-TAE achieves using the full season of data. These results indicate that T3S not only improves in-season classification accuracy and uncertainty quantification but also underscores its practical suitability for operational in-season crop monitoring and decision-support applications, where timely and reliable predictions are critical.
Figure 7: In‐season classification performance at 75th‐ and 50th‐percentile cutoffs, corresponding to end of September and June, respectively, showing (a) accuracy and (b) expected calibration error (ECE), averaged over six folds, for T3S (Ours) and U-TAE.
@@
In many agricultural settings, annotated crop labels are scarce or expensive to obtain, which makes low-data performance a critical requirement for operational viability. To assess robustness under label scarcity, we conducted a six-fold experiment using only 10 % of the training data, randomly sampled once and held constant across methods, and evaluated both T3S and U-TAE on the full test set. As shown in Figure 8, both models incur drops in overall accuracy and increases in ECE when labels are limited, but U-TAE’s performance deteriorates far more rapidly. Remarkably, T3S trained on just 10 % of the data still outperforms U-TAE trained on the full dataset, and its calibration remains substantially more reliable.
Discussion.
The strength of T3S lies in its induced inductive bias: By embedding temperature-informed sampling (the thermal time prior) into the model, we inject domain knowledge that constrains its hypothesis space to solutions that ’make sense’ under known physical or biological rules and prevents overconfident predictions when evidence is weak or conflicting [30, 1]. In practice, the thermal-time prior helps the model concentrate on critical growth stages and discard noise, such as cloud-contaminated observations, steering it toward meaningful phenological signals, shaping more informative learned representations; thus, boosting performance and producing more reliable confidence estimates, particularly in low data regimes, where inductive biases play the greatest role [57, 25]. Consequently, the T3S advantage over the baseline is more pronounced under scarce supervision. It produces well-calibrated uncertainty estimates: clear phenological cues yield high-confidence predictions, while ambiguous or conflicting inputs appropriately result in low confidence.